ZEKE SPECTROSCOPY OF METAL COMPLEXES WITH A MULTIDENTATE LIGAND: M(H_{2}NCH_{2}CH_{2}NH_{2}) (M = Al, Ga, In)

\underline{SHENGGANG LI}, JASON F. FULLER, PARAGRANJITA BHOWMIK, XU WANG, BRADFORD R. SOHNLEIN, and DONG-SHENG YANG, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

Metal coordination to an ethylenediamine ligand may form a chain-like or a cyclic structure. The complex is expected to be a chain if the metal atom is bound to a single nitrogen atom or a five-member ring if the metal is attached to both nitrogen atoms. ZEKE (zero-electron-kinetic-energy) spectra show that M(H_{2}NCH_{2}CH_{2}NH_{2}) and M^{+}(H_{2}NCH_{2}CH_{2}NH_{2}) (M = Al, Ga, In) complexes have a cyclic conformation. These cyclic complexes have significantly lower adiabatic ionization potentials (AIPs) and metal-nitrogen stretching frequencies than simple metal-primary amines. For example, the AIP of Ga(H_{2}NCH_{2}CH_{2}NH_{2}) was determined to be 33322 cm^{−1}, 6008 cm^{−1}less than that of Ga(H_{2}NCH_{3}). The Ga^{+}-N stretching frequency was measured to be 204 cm^{−1}in gallium-ethylenediamine and 299 cm^{−1}in gallium-methylamine.^{a}

aS. Li, G. K. Rothschopf, D. Pillai, B. R. Sohnlein, B. M. Wilson, D. -S. Yang, J. Chem. Phys. 115, 7968 (2001)